In the formula for the mag of the resultant of the difference of 2 vectors 


sq root(a^2 + b^2 - 2abcos$)


is $ the angle between a and b or between a and -b?

Grade:

6 Answers

tejas desai
12 years ago

$ is the angle between A and -B orr.. B and -A

G Narayana Raju
12 years ago

it is the angle between a and -b.

as we know,

we know that a+b=sq.root(a2+b2+2abcosA)

where A is the angle between a and b.

here he asked the difference so we shoud replace b with -b.

then,a-b=sq.root(a2+b2+2a(-b)cosA)

threfore here A=angle between a and-b.

Ashwin Muralidharan IIT Madras
12 years ago

Hi Shayan,

 

Analyse the problem and try to simplify it,

The question says mag of resultant of 2 Vectors = |a - b|

Now |a-b|2 = (a-b).(a-b) = |a|2+|b|2 - 2a.b

or |a-b| = (|a|2+|b|2 - 2a.b)1/2.

So $ is certainly the angle between vector "a" and vector "b".

 

Regards,

Ashwin (IIT Madras).

Ashwin Muralidharan IIT Madras
12 years ago

Shayan,

 

If in the expression you had

(a^2 + b^2 + 2(a)(-b)cos$), then it would have been the angle between a and -b.

 

These kinds of traps are generally set by IITs in the exam, and one needs to be carefule.

In this case it is the angle between a and b only.

 

Regards,

Ashwin (IIT Madras).

Shayan Jaleel
12 years ago

Sorry for being stupid but I really didnt get any of this...

Plz explain again

Ashwin Muralidharan IIT Madras
12 years ago

Hi Shayan,

 

No worries.

The question requires one to find if "$" is the angle between "a" & "b" or "a" & "-b"

So you have to compare the final result with the given expression in the question.

 

You have |a-b| = sqrt{ a^2 + b^2 - 2a.b} as in the above working.

Compare this with the expression in your question.

 

we will have a.b = ab(cos$)

which shows that $ is the abgle between "a" vector and "b" vector.

 

Hope this helps.

 

Best Regards,

Ashwin (IIT Madras).